A particular string resonates in four loops at a frequency of 360Hz .


Give four other frequencies at which it will resonate. Start with the least possible. (F1,F2,F3,F4)
Answer #1

F1=360/4=90 Hz


F2=90*2=180Hz


F3=90*3=270 HZ


F4= 90*5=450 Hz